OpenStudy (lin.ivory):

solve for the horizontal asymptote: lim (√((x^2)+7x-2)) +x x->-∞

OpenStudy (anonymous):

what tools do you have at your disposal?

OpenStudy (lin.ivory):

a scientific calculator?

OpenStudy (anonymous):

lol, i meant what kind of math can you use i think the best bet is to multiply top and bottom by the conjugate of the numerator

OpenStudy (lin.ivory):

its like university calculus? Lol i tried conjugating by multiplying (√((x^2)+7x-2)) -x top and bottom

OpenStudy (anonymous):

the conjugate is \[\sqrt{x^2+7x-2}-x\] when you multiply top and bottom you will get \[\frac{x^2+7x+2-x^2}{\sqrt{x^2+7x-2}-x}\] \[=\frac{7x+2}{\sqrt{x^2+7x+2}-x}\]

OpenStudy (lin.ivory):

yup

OpenStudy (anonymous):

now you want to compute the limit as \(x\to -\infty\)

OpenStudy (lin.ivory):

i tried using the "divide everything by the highest power" method, and i got \[\frac{ 7+\frac{ 2 }{ x } }{ \sqrt{x ^{2}+7x+2}-x }\]

OpenStudy (anonymous):

we can do this with our eyeballs, or do some calculation eyeball method is to say that the numerator is degree 1, and is the denominator because \(\sqrt{x^2} -x\) looks like \(2x\) (since we are going to \(-\infty\) both are positive, then numerator is \(-7x\) so you get \(-\frac{7}{2}\)

OpenStudy (anonymous):

ok we can do it that way too, but you have to divide the denominator by \(x\) as well

OpenStudy (lin.ivory):

so inside the square root, i divide everything by X^2, and outside the square root, i divide the -x by x

OpenStudy (anonymous):

yes but don't forget that you are going to minus infinity so both the radical and(-x\) are positive, but the numerator will be negative that is why it is \(-\frac{7}{2}\)

OpenStudy (lin.ivory):

is there such thing as a -0?

OpenStudy (anonymous):

nope

OpenStudy (lin.ivory):

so 2/(-∞) is still 0

OpenStudy (lin.ivory):

but i dont know what happens at the bottom is it 1-0=1?

OpenStudy (lin.ivory):

why do i divide the (-x) in the denominator by (-x)?

OpenStudy (anonymous):

you don't

OpenStudy (anonymous):

actually, if this is confusing, maybe it is best to divide everything by \(-x\)

OpenStudy (lin.ivory):

ok let me try that

OpenStudy (anonymous):

you have to keep track of the fact that it is the limit as x goes to minus infinity, not plus infinity so both the radical in the denominator, and the term \(-x\) are positive, while the \(7x\) in the numerator is negative

OpenStudy (lin.ivory):

it makes sense by dividing with (-x) but why?

OpenStudy (anonymous):

because of what i wrote above

OpenStudy (anonymous):

i am trying to think of a better way to explain it, but it is alluding me that is why i like the eyeball method better numerator is negative, denominator is positive, numerator looks like \(-7x\) denominator looks like \(2x\)

OpenStudy (lin.ivory):

but why is numerator negative? how do i determine that in other questions?

OpenStudy (anonymous):

or perhaps a better way to explain would be this take the negative of this whole thing, and then take the limit as x goes to \(+\infty\)

OpenStudy (anonymous):

numerator is negative because you are going to minus infinity x goes through numbers like -100,-1000,-1000000000 etc

OpenStudy (lin.ivory):

so every time my limit is approaching x from the negative, i automatically assume the numerator to be negative?

OpenStudy (anonymous):

well not every time, no if it was say \(x^2\) then it would be positive no matter what

OpenStudy (anonymous):

but in this case it is \(7x\) so it is negative for sure as \(x\to -\infty\)

OpenStudy (lin.ivory):

oh yea!! that for sure. But when it is odd exponent, it will automatically be negative?

OpenStudy (anonymous):

yes

OpenStudy (lin.ivory):

THANK YOU SO MUCHHH!! :) that makes a whole lot sense now!

OpenStudy (anonymous):

yw